# STrack: A Reliable Multipath Transport for AI/ML Clusters

### v1

@@ 1. Introduction
Secondly, as AI/ML cluster networks grow to connect tens of thousands or hundreds of thousands nodes, transport’s error resiliency capability is a must in order to achieve high efficiency. Multipathing would make error recovery even more challenging as the transport needs to determine whether a missing packet is a real packet loss or simply a delayed packet due to multipathing. Although RoCEv2’s dependency on lossless Ethernet can help prevent congestion packet drops, bit errors or delayed packets happen often in a vast AI/ML network. RDMA’s go-back-N error recovery mechanism would result in significant performance loss.
@@
This paper presents STrack, the first multipath reliable transport that addresses all aforementioned challenges. STrack employs a novel mechanism that adaptively spray packets to multiple paths without keeping complicated per path state. In addition, we design a window-based congestion control that yields to path selection choice first before cutting down the window in face of pending congestion. We also assume lossy Ethernet as the link layer technology, and design a reliable error recovery mechanism that is based on out-of-order packet counts at receiver NICs to ensure fast packet recovery with minimal spurious retransmission.
Oblivious packet spray distributes packets across multiple paths evenly. Note that the paths here are not the physical paths, but different entropy values, e.g. different UDP source ports, that a sender NIC’s uses to spray traffic. With different entropies in the packet header, ECMP functions at switches would hash them onto different paths. However, oblivious packet spray still experiences hash collisions, and the situation gets worse over time as congested paths accumulate packets as messages continue. In addition, link failures cause full bisectional network becomes asymmetric, which makes oblivious packet spray unbalanced. STrack uses ECN as a congestion signal for the adaptive packet spray algorithm. It keeps a simple bitmap for the entropies that have experienced ECN marks as the congestion state. When an ACK comes back without an ECN mark, that entropy is used to clock out the next new packet. If an ACK comes back with an ECN marked, it is marked in the corresponding position in the bitmap. Next non-marked entropy in a round robin manner is used to clock out new packet. Note that S-track uses only a minimal state, bitmap, to keep congestion information. The bitmap is reset after one or two round trip times.
STrack adopts a sender-only congestion control to handle both fabric congestion and last hop incast. Although a receiver-based scheme can handle a large degree of incast, it is well-known that the AI/ML communication collectives are designed to avoid incast at the application layer. The transport layer incasts do occur but to a much less degree, only up to tens of flows, where sender-based algorithm can easily deal with. In addition, receiver-based congestion control requires sender-based design to combat fabric congestion, also adds more memory footprint to the receiver, which conflicts with the goal of reducing hardware offload complexity.
@@ 2.1. AI/ML Workloads Challenges over RoCEv2
Furthermore, different source of jitters in the system, e.g., stragglers, NIC scheduler or the GPU thread scheduler, can break the order of the collective algorithms, e.g., AlltoAll, and can cause the moderate incast. To demonstrate this point, Figure 3 monitors the queuing delay at all the last-hop switch queues (total of $32$) and traffic arriving rate to each last-hop switch queue as time goes for the AlltoAll experiment with parallel degree of $32$ (A2A/32 in Figure 3) for the RoCEv2, $1$ QPS_PER_CONN setup. Data shows that arrival rates at the last-hop queues exceed $400$ Gbps around $2$ ms, peaking at $489.256$Gbps at $2357$ $\mu$s. This high arrival rate, sustained for approximately $500\mu$s, results in switch queuing delays reaching up to $46\mu$s, equivalent to a queue depth of $2.3$MB. These significant queuing delays underscore the necessity for a more effective congestion control scheme tailored for AI/ML workloads.
Despite the inefficiency of multipath load balancing and congestion control, RoCEv2 also requires network to be lossless due to its inefficient loss recovery mechanism, Go-Back-N. Lossless network usually is achieved by enabling priority flow control (PFC). Prior works have been discussed that PFC storms (flor, ), PFC caused network deadlock (rdmaatscale, ) and head-of-line blocking (irn, ; dcn, ). In addition, given the vast scale of AI/ML clusters, link errors happen all the time. Adopting go-back-N to recover from link bit errors would be very inefficient. To eliminate the lossless network requirement, a selective retransmission mechanism is desired.
@@ 2.2. Hardware Offloaded Transport
Fortunately, as AI/ML workloads do not have a strict requirement for in-order delivery. Similar as SRD (srd, ), STrack provides reliable out-of-order delivery, without the need for a re-order buffer on the NIC. Nonetheless, the new transport must maintain states to support path congestion information for adaptive multipathing, out-of-order packet arrival information for selective acknowledgements and packet retransmissions, and a new window-based congestion control algorithm. Accommodating this per-connection state in an efficient manner to allow a huge number of connections to be maintained in an hardware implementation is a key challenge.
@@
Over the past decade, significant efforts has been made in improving load balancing in data center networks. There are generally two approaches: 1) one is to break down a single flow into multiple subflows, e.g. 4 or 8, and maintains a separate congestion window for each of the subflows, i.e., MPTCP (mptcp, ); 2) the other approach is to spray packets across a large numer of paths, e.g. 128 or 256, and maintaining one congestion window across all paths. The advantages of the subflow approach is that it keeps in-order packet delivery within each subflow. However, due to the per-path congestion state required, this approach limits the number of subflows that can be efficiently maintained in hardware in parallel.
Packets sprayed over a large number of entropies (paths), the workload can be distributed more evenly. The challenge is how keep the path congestion state simple and efficient. With one congestion window governing traffic sent over all paralleled paths, another challenge for multipathing is how to co-ordinate with the congestion control algorithm as various paths could experience different degrees of congestion. Furthermore, even with larger number of entropies, oblivious packet spraying can still perform poorly if the network is asymmetric and the hash collision can not be eliminated completely. STrack adopts a simple and effective method for adaptive packet spraying, which switches the path based on network congestion feedback.
@@
Our first course of action upon a moderate congestion on a few paths, e.g. due to hash collisions, should be shifting some traffic from the over-loaded path to the under-loaded ones. Congestion control should not step in and start cutting the window as reducing the window would limit the number of packets that can be sent to other under-utilized paths. Instead, it diverts some traffic from the over-loaded paths to the under-loaded ones. Congestion control would be necessary during fabric congestion because of over-subscription, network asymmetry, or incast event when multiple senders send to a single receiver.
@@
The design challenge is to decide when to switch paths and when to cut the window. We believe that the earliest congestion signal, indicating onset of congestion, should be used to change path. Without the assumption of network telemetry measures, we make the following key observation: modern switches adopt ECN marking at egress, when a packet exits a queue (dcqcn, ). This particular packet may not have experienced queueing delay, but its ECN mark indicates the queue behind it is building up. This would give us the earliest congestion signal, much faster than RTT or even the change of RTT. Figure 4 shows a simulation of a transient congestion event, when $32$ different sources send traffic to one receiver simultaneously. We measure ECN-marking as well as RTT as time progresses at a sender. As shown in Figure 4, the first received ACK packet for this sender is already ECN-marked while the measured RTT is at base RTT, indicating this particular packet has not experienced any congestion but the queue behind it has built up. The measured RTT or any noticable RTT change does not happen until around 16us, much later after the ECN marked ACK packet has arrived.
With the above observation, we use ECN-based signal for adaptive packet spraying. Only when a majority of paths experienced congestion, we would cut the window. As window adjustment is an quantitative decision, that’s why we take RTT, a multi-bit signal, as the congestion signal for congestion control.
@@ 3. STrack Design
In Scenario #1, the network is not congested, which is the most common network state. Hence, STrack would not interfere as long as the congestion window is at the maximum value, roughly the BDP. If the window is not at the max due to prior congestion episode, STrack increases the congestion window proportionally based on how far the RTT is from the target RTT. If the network is slightly congested, reflected by an marked ECN but with low RTT as in Scenario #2, STrack chooses to switch path while keeping the congestion window intact to allow packets flow to other paths. When RTT becomes large indicating network wide congestion or incast scenarios, STrack cuts window multiplicatively in #3. When RTT is much larger than the target, STrack uses additional congestion signal, achieved Bandwidth Delay Product (BDP), total acknowledged bytes in one base RTT, to quickly convergence in heavy incast scenarios. This allows STrack to converge much faster than the conventional AIMD approach. When congestion goes away, a packet that has experienced long queueing delay will not be ECN marked as no queue built up behind it. In this Scenario #4, STrack actually starts increasing the window additively to avoid starving the link. The ECN marking threshold and the target RTT determine when STrack transitions from Scenario #2 to Scenario #3. The design has a clear separation between the congestion control and multipathing, i.e., the ECN marking threshold is between $25$% and $75$% of Bandwidth Delay Product (BDP) and the target queuing delay is one network RTT, as shown in Figure 7. Initially, we only switch paths based probabilistic ECN marking. As the network congestion increases, the majority of packets are ECN marked and their RTT increases which would trigger congestion window cut. Note that Figure 7 reflects the symbolic queueing view from congestion control’s angle while multiple paths’ queue are involved in a real network.
STrack assumes that switches drop packets silently when no buffer space is available, i.e., trimming (ndp, ; cp, ) or Back-to-sender(sfc, ; bolt, ) is not present. Furthermore, STrack’s reliability mechanisms must also align with coalescing ACKs (smith2010tcp, ; johnson2018tcp, ) to accommodate the reality that packet processing rates grow at a slower pace than link speeds. STrack provides a packet loss detection mechanism (Section 3.3) across three different time scales: (1) out-of-order packets counters; (2) probing-based approach; (3) timeout. Once loss is detected, the retransmited packets are sent if the congestion window is allowed.
@@
1:procedure on_receiving_ack
2:    ecn = ack.ecn
3:    path_id = ack.path_id
4:    update_ecn_bitmap(ecn, path_id)
5:    if measured_rtt ¡ base_rtt then
6:         base_rtt = measured_delay
7:    end if
8:    measured_Qdelay = measured_rtt - base_rtt
9:    probe_timer_ts = now + 3* net_base_rtt
10:    if ack_for_probe $and$ delay $<$ 22*2 ∗ net_base_rtt $and$ achievedBDP == 0 then
11:         retransmit_packets()
12:    end if
13:    achievedBDP = update_achievedBDP(ack, now)
14:    cwnd = adjust_cwnd(ecn, measured_Qdelay, achievedBDP)
15:    loss_detection(ack)
16:end procedure
Algorithm 2 STrack AchievedBDP
1:procedure update_achievedBDP(ack, now)
3:    rx_count += ack_for_probe? 0, acked_bytes
4:    if can_clear_byte then
5:         achievedBDP = rx_count
6:         rx_count = 0
7:         rxcount_clear_ts = now
8:    end if
9:end procedure
Algo 1 shows the overall-design of STrack. We assume an ACK packet echoes back the ecn mark, path_id and timestamp to the sender. Upon receiving an ACK, STrack checks: 1) whether an ECN is marked or not; 2) measured queuing delay (Line #8); 3) Received ACKed bytes (Line #13), which are our congestion signals. STrack updates the ECN bitmap with the echoed path_id and ecn signal (Line #4), which are used by multipath load balancing algorithm (in Algo 3). The main congestion control algorithm adjust_cwnd is specified in Algo 4, which we describe separately later. The packet loss detection algorithm is described in Section 3.3. Upon sending out a packet, we choose the right path id (Algo 3), which is covered later.
@@ 3.1. Adaptive Load Balancing
Algo 3 explains STrack adaptive load balancing algorithm. It initializes a bitmap with all the paths available, i.e., set bit as 00 and set the $next\_path\_id$ as invalid, i.e., a negative value.
Upon receiving an ACK packet, STrack remembers the entropy value that is returned ecn-free in next_path_id and label the path_id as good path in the bitmap (Line #7-8 in Algo 3). When sending data packets, its uses entropy values in a round-robin fashion until an ACK is received. If the ACK is not ECN-marked, its entropy would be used for the packet that is being sent (Line #13 in Algo 3). Otherwise, it checks the bitmap in a round-robin fashion. The total paths are dynamically changed with the congestion window (Line #17 in Algo 3). This is because the path congestion information is fresh within a window. For example, if we maintain a bitmap size of $256$, while the congestion window is one, it requires $256$ RTTs to update the whole bitmap, and the congestion information is already outdated when the round-robin turns to a particular entropy. While checking the bitmap, if the path is ecn-marked, we skip the path and clear the bit for the first skipped path (Line #23 in Algo 3).
@@ 3.2. Congestion Control
$\beta$ Used for additive increase 55*5 ∗bdp_sf
$\eta$ Used for fairness $0.15$*bdp_sf
$\alpha$ Used for the RTT gain increase $4.0$*bdp_sf* delay_sf/base_rtt
@@
2:    update_achievedBDP(ack, now)
3:    can_decrease = now - last_decrease_ts $>$ base_rtt
4:    can_fairness_shuffle = now - last_selfai_ts $>$ base_rtt
5:    avg_delay = avg_delay*(1 - $ewma$) +$ewma$*delay
6:    if !ecn and delay $>$ $target\_Qhigh$  then
7:         cwnd $\leftarrow$ cwnd + $\frac{\beta}{cwnd}$
8:    else if !ecn and delay $<$ target_Qdelay  then
9:         cwnd $\leftarrow$ cwnd + $\alpha$*(target_Qdelay - delay)/cwnd
10:    else if can_decrease and avg_delay ¿ target_Qdelay  then
11:         if delay $>$ $target\_Qhigh$ and achievedBDP ¡ max_cwnd/8 then
12:             cwnd = achievedBDP
13:             last_decrease_ts = now
14:         else if delay ¿ target_Qdelay then
15:             cwnd = cwnd* max( 1 - $\gamma$ (avg_delay - target_Qdelay)/avg_delay, $0.5$)
16:             last_decrease_ts = now
@@
18:    end if
19:    if can_fairness_shuffle then
20:         cwnd = cwnd + $\eta$
21:         last_selfai_ts = now
22:    end if
23:    return cwnd
24:end procedure
Algo 4 shows the design of STrack’s congestion control algorithm. Table 1 lists the configuration parameters for the performance tuning and use $bdp\_sf$ and $delay\_sf$ as BDP scaling factor and delay scaling factor to adapt to various network speeds and network latency. $target\_Qdelay$ is target queuing delay that we expect the congestion control to be stabilized at. The algorithm maintains one congestion window to handle all the paths congestion.
For efficiency, STrack increases the rate in two different situations. First, when ecn is not marked and current measured queuing delay (delay in Algo 4) is below the target_Qdelay, the rate increase is proportional to the difference between current RTT and the target_delay (Line #9). Secondly, we increase the rate by a constant value when ecn is not marked but measured queuing delay is twice the target delay (Line #7). This is a situation when this packet incurred a long delay, but the queue length has come down drastically (ecn is not marked), we should increase the rate to avoid link starvation. There are two cases where we reduce the window: depending on the current measured queuing delay is greater than $target\_Qhigh$ or not. If it is greater, we use the achieved BDP (achievedBDP in Algo 4) as our window to quickly converge to a lower rate (Line #12). Otherwise, we do multiplicative decrease based on how much average measured queuing delay (avg_delay in Algo 4) is over the target_Qdelay (Line #15). To ensure fast convergence of fairness, we periodically add a constant small increase to the congestion window (Line #20). Note that this addition would assist the algorithm to have slightly more packets to explore different paths.

### v2

@@ 1. Introduction
Secondly, as AI/ML cluster networks grow to connect tens of thousands or hundreds of thousands nodes, transport’s error resiliency capability is a must in order to achieve high efficiency. Multipathing would make error recovery even more challenging as the transport needs to determine whether a missing packet is a real packet loss or simply a delayed packet due to multipathing. Although RoCEv2’s dependency on lossless Ethernet can help prevent congestion packet drops, bit errors or packet delays happen often in a vast AI/ML network. RDMA’s go-back-N error recovery mechanism would result in significant performance loss.
@@
This paper presents STrack, the first reliable multipath transport that addresses all aforementioned challenges. STrack employs a novel mechanism that adaptively spray packets to multiple paths without keeping complicated per path state. In addition, we design a window-based congestion control that yields to path selection choice first before cutting down the window in face of pending congestion. We also assume lossy Ethernet as the link layer technology, and design a reliable error recovery mechanism that is based on out-of-order packet counts at receiver NICs to ensure fast packet recovery with minimal spurious retransmission.
STrack uses egress-marked ECN as a congestion signal for the adaptive packet spray algorithm instead of oblivious packet spray that distributes packets across multiple paths evenly. Note that the paths here are not the physical paths, but different entropy values, e.g. different UDP source ports, that a sender NIC’s uses to spray traffic. With different entropies in the packet header, ECMP functions at switches would hash them onto different paths. However, oblivious packet spray still experiences hash collisions, and the situation gets worse over time as congested paths accumulate packets as messages continue. In addition, link failures cause full bisectional network becomes asymmetric, which makes oblivious packet spray unbalanced. STrack keeps a simple bitmap for the entropies that have experienced ECN marks as the congestion state. When an ACK comes back without an ECN mark, that entropy is used to clock out the next new packet. If an ACK comes back with an ECN marked, it is marked in the corresponding position in the bitmap. Next non-marked entropy in a round robin manner is used to clock out new packet. Note that S-track uses only a minimal state, bitmap, to keep congestion information. The bitmap is reset after one or two round trip times.
STrack adopts a sender-only congestion control to handle both fabric congestion and last hop incast. Although a receiver-based scheme can handle a large degree of incast, it is well-known that the AI/ML communication collectives are designed to avoid incast at the application layer. The transport layer incasts do occur but to a much less degree, only up to tens of flows, where sender-based algorithm can easily deal with. In addition, receiver-based congestion control requires sender-based design to combat fabric congestion, and adds more memory footprint to the receiver, which conflicts with our goal of reducing hardware offload complexity.
@@ 2.1. AI/ML Workloads Challenges over RoCEv2
Furthermore, different source of jitters in the system, e.g., stragglers, NIC scheduler or the GPU thread scheduler, can break the order of the collective algorithms, e.g., AlltoAll, and can cause the moderate incast. To demonstrate this point, Figure 3 monitors the queuing delay at all the last-hop switch queues (total of $32$ queues) and traffic arriving rate to each last-hop switch queue as time goes for the AlltoAll experiment with parallel degree of $32$ (A2A/32 in Figure 3) for the RoCEv2, $1$ QPS_PER_CONN setup. Data shows that arrival rates at the last-hop queues exceed $400$ Gbps on multiple ocasions. For example, around $2$ ms, the arrival rate peaks at $489.256$Gbps at $2357$ $\mu$s. This high arrival rate, sustained for approximately $500\mu$s, results in switch queuing delays reaching up to $46\mu$s, equivalent to a queue depth of $2.3$MB. These significant queuing delays underscore the necessity for a more effective congestion control scheme tailored for AI/ML workloads.
In addition to the issues of single path load balancing and rate-based congestion control, RoCEv2 also requires network to be lossless due to its inefficient loss recovery mechanism, Go-Back-N. Lossless network usually is achieved by enabling priority flow control (PFC). Prior works have been discussed that PFC storms (flor, ), PFC caused network deadlock (rdmaatscale, ) and head-of-line blocking (irn, ; dcn, ). In addition, given the vast scale of AI/ML clusters, link errors happen all the time. Adopting go-back-N to recover from link bit errors would be very inefficient. To eliminate the lossless network requirement, a selective retransmission mechanism is desired.
@@ 2.2. Hardware Offloaded Transport
Fortunately, as AI/ML workloads do not have a strict requirement for in-order delivery. Similar as SRD (srd, ), STrack provides reliable out-of-order delivery, without the need for a re-order buffer on the NIC. Nonetheless, the new transport must maintain states to support path congestion information for adaptive multipathing, out-of-order packet arrival information for selective acknowledgements and packet retransmissionson top of a new window-based congestion control algorithm. Accommodating this per-connection state in an efficient manner to allow a huge number of connections to be maintained in an hardware implementation is a key challenge.
@@
Over the past decade, significant efforts has been made in improving load balancing in data center networks. There are generally two approaches: 1) one is to break down a single flow into multiple subflows, e.g. 4 or 8, and maintains a separate congestion window for each of the subflows, i.e., MPTCP (mptcp, ); 2) the other approach is to spray packets across a large numer of paths, e.g. 128 or 256, and maintaining one congestion window across all paths (srd, ). The advantages of the subflow approach is that it keeps in-order packet delivery within each subflow. However, due to the per-path congestion state required, this approach limits the number of subflows that can be efficiently maintained in hardware in parallel.
If packets are sprayed over a large number of entropies (paths), the workload can be distributed more evenly. However, even with larger number of entropies, oblivious packet spraying can still perform poorly if the network is asymmetric and the hash collision can not be eliminated completely. Adaptive packet spray takes into account paths’ congestion. The challenge is how to keep the path congestion state simple and efficient. If one congestion window governing traffic sent over all paralleled paths, the additional challenge is how multipathing co-ordinates with the congestion control algorithm as various paths could experience different degrees of congestion.
@@
Ideally, the first course of action upon a moderate congestion on a few paths should be shifting some traffic from the over-loaded path to the under-loaded ones. Congestion control should not step in and start cutting the window as reducing the window would limit the number of packets that can be sent to other under-utilized paths. Instead, it diverts some traffic from the over-loaded paths to the under-loaded ones. Congestion control would be necessary during fabric congestion because of over-subscription, network asymmetry, or incast event when multiple senders send to a single receiver.
@@
The design challenge is deciding when to switch paths and when to cut the window. We believe that the earliest congestion signal, indicating onset of congestion, should be used to change path. Without the assumption of network telemetry measures, we make the following key observation: modern switches adopt ECN marking at egress, when a packet exits a queue (dcqcn, ). This particular packet may not have experienced queueing delay, but its ECN mark indicates the queue behind it is building up. This would give us the earliest congestion signal, much faster than RTT or even the change of RTT. Figure 4 shows a simulation of a transient congestion event, when $32$ different sources send traffic to one receiver simultaneously. We measure ECN-marking as well as RTT as time progresses at a sender. As shown in Figure 4, the first received ACK packet for this sender is already ECN-marked while the measured RTT is at base RTT, indicating this particular packet has not experienced any congestion but the queue behind it has built up. The measured RTT or any noticable RTT change does not happen until around 16us, much later after the ECN marked ACK packet has arrived.
With the above observation, we believe that ECN is the most timely signal for adaptive packet spraying. Only when a majority of paths experienced congestion, we would cut the window. As window adjustment is an quantitative decision, that’s why we take RTT, a multi-bit signal, as the congestion signal for congestion control.
@@ 3. STrack Design
In Scenario #1, the network is not congested, which is the most common network state. Hence, STrack would not interfere as long as the congestion window is at the maximum value, roughly the BDP. If the window is not at the max due to prior congestion episode, STrack increases the congestion window proportionally based on how far the RTT is from the target RTT. If the network is slightly congested, reflected by an marked ECN but with low RTT as in Scenario #2, STrack chooses to switch path while keeping the congestion window intact to allow packets flow to other paths. When RTT becomes large indicating network wide congestion or incast scenarios, STrack cuts window multiplicatively in #3. When RTT is much larger than the target, STrack uses additional congestion signal, achieved Bandwidth Delay Product (BDP), total acknowledged bytes in one base RTT, to quickly reaches the operating point in heavy incast scenarios. This allows STrack to converge much faster than the conventional AIMD approach. When congestion goes away, a packet that has experienced long queueing delay will not be ECN marked as no queue built up behind it. In this Scenario #4, STrack actually starts increasing the window additively to avoid starving the link. The ECN marking threshold and the target RTT determine when STrack transitions from Scenario #2 to Scenario #3.
The design has a clear separation between the congestion control and multipathing, i.e., the ECN marking threshold is between $25$% and $75$% of Bandwidth Delay Product (BDP) and the target queuing delay is one network RTT, as shown in Figure 7. Initially, we only switch paths based probabilistic ECN marking. As the network congestion increases, the majority of packets are ECN marked and their RTT increases which would trigger congestion window cut. The adaptive packet spray is still on-going even if with a window cut. Note that Figure 7 reflects the symbolic queueing view from congestion control’s angle while multiple paths’ queue are involved in a real network.
STrack assumes that switches drop packets silently when no buffer space is available, i.e., trimming (ndp, ; cp, ) or Back-to-sender(sfc, ; bolt, ) is not present. Furthermore, STrack’s reliability mechanisms must also align with coalescing ACKs (smith2010tcp, ; johnson2018tcp, ) to accommodate the reality that packet processing rates grow at a slower pace than link speeds. STrack provides a packet loss detection mechanism (Section 3.3): (1) out-of-order packets counters; (2) probing-based approach; (3) timeout. Once loss is detected, the retransmited packets are sent if the congestion window is allowed.
@@
1:procedure on_sending_packet(packet)
2:    packet.path_id = choose_path()
3:end procedure
4:procedure on_receiving_ack
5:    ecn = ack.ecn
6:    path_id = ack.path_id
7:    update_ecn_bitmap(ecn, path_id)
8:    if measured_rtt ¡ base_rtt then
9:         base_rtt = measured_delay
10:    end if
11:    measured_Qdelay = measured_rtt - base_rtt
12:    probe_timer_ts = now + 3* net_base_rtt
13:    if ack_for_probe $and$ delay $<$ 22*2 ∗ net_base_rtt $and$ achievedBDP == 0 then
14:         retransmit_packets()
15:    end if
16:    achievedBDP = update_achievedBDP(ack, now)
17:    cwnd = adjust_cwnd(ecn, measured_Qdelay, achievedBDP)
18:    loss_detection(ack)
19:end procedure
Algo 1 shows the overall-design of STrack. We assume an ACK packet echoes back the ecn mark, path_id and timestamp to the sender. Upon receiving an ACK, STrack checks: 1) whether an ECN is marked or not; 2) measured queuing delay (Line #8); 3) Received ACKed bytes (Line #13), which are our congestion signals. STrack updates the ECN bitmap with the echoed path_id and ecn signal (Line #4), which are used by multipath load balancing algorithm (in Algo 2). The main congestion control algorithm adjust_cwnd is specified in Algo 3, which we describe separately later. The packet loss detection algorithm is described in Section 3.3. Upon sending out a packet, we choose the right path id (Algo 2), which is covered later.
@@ 3.1. Adaptive Load Balancing
Algo 2 explains STrack adaptive load balancing algorithm. It initializes a bitmap with all the paths available, i.e., set bit as 00 and set the $next\_path\_id$ as invalid, i.e., a negative value.
Upon receiving an ACK packet, STrack remembers the entropy value that is returned ecn-free in next_path_id and label the path_id as good path in the bitmap (Line #7-8 in Algo 2). When sending data packets, its initially uses entropy values in a round-robin fashion until an ACK is received. If the ACK is not ECN-marked, its entropy would be used for the packet that is being sent (Line #13 in Algo 2). Otherwise, it checks the bitmap in a round-robin fashion. The total paths are dynamically changed with the congestion window (Line #17 in Algo 2). This is because the path congestion information is fresh within a window. For example, if we maintain a bitmap size of $256$, while the congestion window is one, it requires $256$ RTTs to update the whole bitmap, and the congestion information is already outdated when the round-robin turns to a particular entropy. While checking the bitmap, if the path is ecn-marked, we skip the path and clear the bit for the first skipped path (Line #23 in Algo 2).
@@ 3.2. Congestion Control
$\beta$ Used for additive increase 55*5 ∗MTU*bdp_sf
$\eta$ Used for fairness $0.15$*MTU*bdp_sf
$\alpha$ Used for the RTT gain increase $4.0$*bdp_sf* delay_sf*MTU/base_rtt
@@
2:    can_decrease = now - last_decrease_ts $>$ base_rtt
3:    can_fairness_shuffle = now - last_selfai_ts $>$ base_rtt
4:    avg_delay = avg_delay*(1 - $ewma$) +$ewma$*delay
5:    if !ecn and delay $>$ $target\_Qhigh$  then
6:         cwnd $\leftarrow$ cwnd + $\frac{\beta}{cwnd}$
7:    else if !ecn and delay $<$ target_Qdelay  then
8:         cwnd $\leftarrow$ cwnd + $\alpha$*(target_Qdelay - delay)/cwnd
9:    else if can_decrease and avg_delay ¿ target_Qdelay  then
10:         if delay $>$ $target\_Qhigh$ and achievedBDP ¡ max_cwnd/8 then
11:             cwnd = achievedBDP
12:             last_decrease_ts = now
13:         else if delay ¿ target_Qdelay then
14:             cwnd = cwnd* max( 1 - $\gamma$ (avg_delay - target_Qdelay)/avg_delay, $0.5$)
15:             last_decrease_ts = now
16:         end if
@@
18:    if can_fairness_shuffle then
19:         cwnd = cwnd + $\eta$
20:         last_selfai_ts = now
21:    end if
22:    return cwnd
23:end procedure
Algorithm 4 STrack AchievedBDP
1:procedure update_achievedBDP(ack, now)
3:    rx_count += ack_for_probe? 0, acked_bytes
4:    if can_clear_byte then
5:         achievedBDP = rx_count
6:         rx_count = 0
7:         rxcount_clear_ts = now
8:    end if
9:end procedure
Algo 3 shows the design of STrack’s congestion control algorithm. Table 1 lists the configuration parameters for the performance tuning and use $bdp\_sf$ and $delay\_sf$ as BDP scaling factor and delay scaling factor to adapt to various network speeds and network latency. $target\_Qdelay$ is target queuing delay that we expect the congestion control to be stabilized at. The algorithm maintains one congestion window to handle all the paths congestion.
For efficiency, STrack increases the rate in two different situations. First, when ecn is not marked and current measured queuing delay (delay in Algo 3) is below the target_Qdelay. This happens when a congestion episode has just passed, the algorithm increases the rate that is proportional to the difference between current RTT and the target_delay (Line #8). Secondly, we increase the rate by a constant value when ecn is not marked but measured queuing delay is twice the target delay (Line #6). This is a situation when this packet incurred a long delay, but the queue length has come down drastically (ecn is not marked), we should increase the rate to avoid link starvation. There are two cases where we reduce the window: depending on the current measured queuing delay is greater than $target\_Qhigh$ or not. If it is greater, we use the achieved BDP (achievedBDP in Algo 4) as our window to quickly converge to a lower rate (Line #11). Otherwise, we do multiplicative decrease based on how much average measured queuing delay (avg_delay in Algo 3) is over the target_Qdelay (Line #14). To ensure fast convergence of fairness, we periodically add a constant small increase to the congestion window (Line #19). Note that this addition would assist the algorithm to have a slightly bigger window in order for packets to explore different paths.